# Single-BS Simultaneous Environment Sensing and UE Localization without LoS Path by Exploiting Near-Field Scatterers

### v1

@@ III-C Delay Estimation
$\left(\bm{F}_{\mathrm{Rx},l}\right)_{n,\gamma}$ $=\underbrace{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l}b_{n,\gamma}e^{j2\pi(\gamma f_{D,l}T_{s}-n\Delta f\tau_{s,l})}}_{\mathrm{desired\,signal}}$ (15)
$+\underbrace{\sum_{l^{\prime}\neq l}^{L}{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l^{\prime}}b_{n,\gamma}e^{j2\pi(\gamma f_{D,l^{\prime}}T_{s}-n\Delta f\tau_{s,l^{\prime}})}}}_{\mathrm{inter-path\,interference}}\!+\left(\!\bm{Z}_{l}\right)_{n,\gamma},$
@@
$\left(\bm{F}_{l}\right)_{n,\gamma}$ $=\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l}e^{j2\pi(\gamma f_{D,l}T_{s}-n\Delta f\tau_{s,l})}$ (16)
$+\sum_{l^{\prime}\neq l}^{L}{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l^{\prime}}e^{j2\pi(\gamma f_{D,l^{\prime}}T_{s}-n\Delta f\tau_{s,l^{\prime}})}}+(\tilde{\bm{Z}}_{l})_{n,\gamma},$

### v3

@@ III-C Delay Estimation
$\left(\bm{F}_{\mathrm{Rx},l}\right)_{n,\gamma}$ $=\underbrace{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l}b_{n,\gamma}e^{j2\pi(\gamma f_{D,l}T_{O}-n\Delta f\tau_{s,l})}}_{\mathrm{desired\,signal}}$ (15)
$+\underbrace{\sum_{l^{\prime}\neq l}^{L}{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l^{\prime}}b_{n,\gamma}e^{j2\pi(\gamma f_{D,l^{\prime}}T_{O}-n\Delta f\tau_{s,l^{\prime}})}}}_{\mathrm{inter-path\,interference}}\!+\left(\!\bm{Z}_{l}\right)_{n,\gamma},$
@@
$\left(\bm{F}_{l}\right)_{n,\gamma}$ $=\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l}e^{j2\pi(\gamma f_{D,l}T_{O}-n\Delta f\tau_{s,l})}$ (16)
$+\sum_{l^{\prime}\neq l}^{L}{\!\hat{\bm{f}}_{l}^{H}\bm{h}_{l^{\prime}}e^{j2\pi(\gamma f_{D,l^{\prime}}T_{O}-n\Delta f\tau_{s,l^{\prime}})}}+(\tilde{\bm{Z}}_{l})_{n,\gamma},$